For a lab report, when calculating a concentration using this formula, Concentration (g/L) = (# drops of dye added to well)(0.030 g/L) / (10 drops) which would be the correct way to answer in sig figs?

Concentration of Well 1 (g/L) = (1)(0.030 g/L)/(10 drops) = .003

Concentration of Well 2 (g/L) = (2)(0.030 g/L)/(10 drops) = .006

Concentration of Well 3 (g/L) = (3)(0.030 g/L)/(10 drops) = .009

Concentration of Well 4 (g/L) = (4)(0.030 g/L)/(10 drops) = .012 or .01 ??

Concentration of Well 5 (g/L) = (5)(0.030 g/L)/(10 drops) = .015 or .01 ??

Concentration of Well 6 (g/L) = (6)(0.030 g/L)/(10 drops) = .018 or .01 ??

Concentration of Well 7 (g/L) = (1)(0.030 g/L)/(10 drops) = .021 or .02 ??

Concentration of Well 8 (g/L) = (8)(0.030 g/L)/(10 drops) = .024 or .02 ??

Concentration of Well 9 (g/L) = (9)(0.030 g/L)/(10 drops) = .027 or .02 ??

Concentration of Well 10 (g/L) = (10)(0.030 g/L)/(10 drops) = .03
Answer #1

there are three rules on determining how many significant figure in a particular number

significant figure : The number of digits used to express any decimal or real number a measured or calculated quantity. If we express a number beyond the place to which we have actually measured

from question ::

1) Concentration of Well 1 (g/L) = (1)(0.030 g/L)/(10 drops) = .003  

there   significant figure is =3

2) Concentration of Well 2 (g/L) = (2)(0.030 g/L)/(10 drops) = .006

   there   significant figure is =3

3)Concentration of Well 3 (g/L) = (3)(0.030 g/L)/(10 drops) = .009

there   significant figure is =3

4)Concentration of Well 4 (g/L) = (4)(0.030 g/L)/(10 drops) = .012 or .01 ??

well4 is =0.12 is correct way

there   significant figure is =2

5)Concentration of Well 5 (g/L) = (5)(0.030 g/L)/(10 drops) = .015 or .01 ??

well 5 is correct way = .015 or its equal to .02 not 0.01

so correct way is = .015

there   significant figure is =3

6)Concentration of Well 6 (g/L) = (6)(0.030 g/L)/(10 drops) = .018 or .01 ??

well 5 is correct way = .018 or its equal to .02 not 0.01

so correct way is = .018

there   significant figure is =3

7)Concentration of Well 7 (g/L) = (1)(0.030 g/L)/(10 drops) = .021 or .02 ??

there both are correct but .021 better than .02

there   significant figure is of ,021=3

8)Concentration of Well 8 (g/L) = (8)(0.030 g/L)/(10 drops) = .024 or .02 ??

there both are correct but .024 better than .02

there   significant figure is .024 is=3

9)Concentration of Well 9 (g/L) = (9)(0.030 g/L)/(10 drops) = .0 27 equals to .3

but correct way is =.027

there   significant figure is =3

10)Concentration of Well 10 (g/L) = (10)(0.030 g/L)/(10 drops) = .03

there   significant figure is =2